A die is thrown and a card is drown from well shuffled deck of 52 playing card at the same time , what is the probability of that an out come consist of
A) even number from the die and diamond from the card
B) a2 from the die and a pictured card from playing card
C) a prime number from the die and a king from the card
D) a composite number from the die and a red card from the playing card

Answer by Boreal(15235)

A. (1/2)*(1/4)=1/8
B. (1/6)*(3/13)=1/26
C. (1/2)*(1/13)=1/26; the prime numbers are 2,3,5
D. (1/3)*(1/2)=1/6